# Generating Images with 3D Annotations Using Diffusion Models

### v3

@@ 3.1 Background: Diffusion Models with Text and 2D Visual Prompts
Algorithm 1 Generating images using our 3D-DST 1:  Input: 3D shape repository $\mathcal{C}$, $w$, $k$}. 2:  Output: Images with 3D annotations {$I_{\text{final}},y_{\text{3D}}$}. 3:  for iterations do 4:    Get a CAD model $\mathcal{C}$ from the 3D shape repository; 5:    Load the class name $w$ and keyword $k$ of $\mathcal{C}$; 6:    Load the camera intrinsic matrix $K$; 7:    Generate randomized camera extrinsic matrices $\{\xi\}$; 8:    for different camera extrinsic matrix $\xi$ do 9:      // 3D visual prompt generation 10:      Render sketch image $I_{\text{render}}$ using $\xi$ and $K$ by Eq. 4; 11:      Produce 3D annotations $y_{\text{3D}}$ of $I_{\text{render}}$; 12:      Compute edge map $\mathcal{E}_{\text{3D}}$ using Canny Edge by Eq. 5 ; 13:       // Text prompt generation 14:       Create text prompts $\mathcal{T}$ using $w$ and $k$ by Eq. 6; 15:       // Image generation 16:      Initialize a random noise $z_{T}$; 17:      for $t$ in $T,\cdots,1$ do 18:        Denoise $z_{t}$ to $z_{t-1}$ using $\mathcal{T}$ and $\mathcal{E}_{\text{3D}}$ by Eq. 3; 19:      Generate images $I_{\text{final}}=\mathcal{D}(z_{0})$. 20:  Collect the generated dataset {$I_{\text{final}},y_{\text{3D}}$}.
@@ 4 Experiments
  w / 3D-DST (ours) 87.36 (normal-$\uparrow$3.52) 88.96 (normal-$\uparrow$4.40) 88.11 (normal-$\uparrow$3.83) 92.18 (normal-$\uparrow$0.84) 92.44 (normal-$\uparrow$0.94) 91.50 (normal-$\uparrow$0.72) 90.47 (normal-$\uparrow$1.78)
@@
  w / 3D-DST (ours) 56.12 (normal-$\uparrow$6.16) 56.65 (normal-$\uparrow$7.04) 56.74 (normal-$\uparrow$6.21) 69.69 (normal-$\uparrow$2.50) 69.21 (normal-$\uparrow$2.81) 59.97 (normal-$\uparrow$4.98) 68.42 (normal-$\uparrow$3.24)
@@ Appendix F Analyzing the Quality of the Automatic 3D annotations
  w / 3D-DST (ours) 88.96 (normal-$\uparrow$4.40) 92.18 (normal-$\uparrow$0.84) 56.65 (normal-$\uparrow$7.04) 69.69 (normal-$\uparrow$2.50)
@@
  w / DST-3D (ours) 42.34 (normal-$\uparrow$0.84) 42.74 (normal-$\uparrow$1.09)
  w / DST-3D + camera aug (ours) 42.86 (normal-$\uparrow$1.36) 43.19 (normal-$\uparrow$1.54)

### v4

@@ 3.1 Background: Diffusion Models with Text and 2D Visual Prompts
Algorithm 1 Generating images using our 3D-DST
1:  Input: 3D shape repository $\mathcal{C}$, $w$, $k$}.
2:  Output: Images with 3D annotations {$I_{\text{final}},y_{\text{3D}}$}.
3:  for iterations do
4:    Get a CAD model $\mathcal{C}$ from the 3D shape repository;
5:    Load the class name $w$ and keyword $k$ of $\mathcal{C}$;
6:    Load the camera intrinsic matrix $K$;
7:    Generate randomized camera extrinsic matrices $\{\xi\}$;
8:    for different camera extrinsic matrix $\xi$ do
9:      // 3D visual prompt generation
10:      Render sketch image $I_{\text{render}}$ using $\xi$ and $K$ by Eq. 4;
11:      Produce 3D annotations $y_{\text{3D}}$ of $I_{\text{render}}$;
12:      Compute edge map $\mathcal{E}_{\text{3D}}$ using Canny Edge by Eq. 5 ;
13:       // Text prompt generation
14:       Create text prompts $\mathcal{T}$ using $w$ and $k$ by Eq. 6;
15:       // Image generation
16:      Initialize a random noise $z_{T}$;
17:      for $t$ in $T,\cdots,1$ do
18:        Denoise $z_{t}$ to $z_{t-1}$ using $\mathcal{T}$ and $\mathcal{E}_{\text{3D}}$ by Eq. 3;
19:      Generate images $I_{\text{final}}=\mathcal{D}(z_{0})$.
20:  Collect the generated dataset {$I_{\text{final}},y_{\text{3D}}$}.
@@ 4 Experiments
  w / 3D-DST (ours) 87.36 ($\uparrow$3.52) 88.96 ($\uparrow$4.40) 88.11 ($\uparrow$3.83) 92.18 ($\uparrow$0.84) 92.44 ($\uparrow$0.94) 91.50 ($\uparrow$0.72) 90.47 ($\uparrow$1.78)
@@
  w / 3D-DST (ours) 56.12 ($\uparrow$6.16) 56.65 ($\uparrow$7.04) 56.74 ($\uparrow$6.21) 69.69 ($\uparrow$2.50) 69.21 ($\uparrow$2.81) 59.97 ($\uparrow$4.98) 68.42 ($\uparrow$3.24)
@@ Appendix F Analyzing the Quality of the Automatic 3D annotations
  w / 3D-DST (ours) 88.96 ($\uparrow$4.40) 92.18 ($\uparrow$0.84) 56.65 ($\uparrow$7.04) 69.69 ($\uparrow$2.50)
@@
  w / DST-3D (ours) 42.34 ($\uparrow$0.84) 42.74 ($\uparrow$1.09)
  w / DST-3D + camera aug (ours) 42.86 ($\uparrow$1.36) 43.19 ($\uparrow$1.54)